Locate the centroid of the volume bounded by the equation y^2 = 4x, x = 1 and the x-axis and revolving about the x-axis.


1
Expert's answer
2022-04-13T09:16:30-0400



The equations for calculating the centroidal coordinates of a volume of arbitrary shape are given as

"\overline{x}=\frac{\displaystyle\int_V xdV}{\displaystyle\int_V dV}=\frac{\displaystyle\int_V xdV}{\displaystyle V}"

"\overline{y}=\frac{\displaystyle\int_V ydV}{\displaystyle\int_V dV}=\frac{\displaystyle\int_V ydV}{\displaystyle V}"

"\overline{z}=\frac{\displaystyle\int_V zdV}{\displaystyle\int_V dV}=\frac{\displaystyle\int_V zdV}{\displaystyle V}"

Let's find "\overline{x}":

"y^2=4x" , "x=1"

"y=\pm2\sqrt x" , "0\le x\le1"

"dV=\pi y^2(x)dx=\pi (2\sqrt x)^2dx=\pi\cdot4xdx"

"\displaystyle\int_V xdV=\displaystyle\int_0^1 x\pi\cdot4xdx=\frac43\pi x^3|_0^1=\frac43\pi"

Volume can be found using the formula:

"V=\displaystyle\int_V dV=\pi\displaystyle\int_0^1y^2(x)dx=\pi\displaystyle\int_0^1(2\sqrt x)^2dx""=\pi\displaystyle\int_0^14xdx=\pi\cdot2x^2|_0^1=2\pi"

"\overline{x}=\frac{\displaystyle\int_V xdV}{\displaystyle\int_V dV}=\displaystyle\frac43\pi/(2\pi)=\frac23"

No necessary to calculate "\overline{y}" and "\overline{z}" . Since the body is obtained by rotation around the x axis, it means that it is symmetrical along the x axis and the centroid point will lie on the x axis. Therefore we can conclude

"\overline{y}=0"

"\overline{z}=0"

Answer: Coordinates of the centroid: ("\frac23,0,0)" .